# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ I-B7 Release 19: 5G-Advanced Evolution
Release 19 (1/2024-6/2025) will be the first evolution of 5G-Advanced. 3GPP agreed on its content in December 2023, and the work began in the beginning of 2024. Release 19 continues the 5G evolution toward 6G, and can be seen largely as preparation studies toward 6G, while still continuing the development based on commercial deployment needs. One focus will be on developing further use cases, such as the NTN evolution for NR and IoT, femto cells, wireless access backhaul, low-power wake-up signals, ambient IoT, and multi-hop relays. Another focus area is to further improve the eMBB experience, such as extended reality (XR), mobility, MIMO operation, and duplex. Some interesting starting studies toward 6G technologies include channel modeling for 7-24GHz, which is one of the important new 6G bands in the 6G roadmap and channel modeling for sensing use cases. In addition, improvements of the other key performance indicators will continue in terms of network energy saving enhancements, AI/ML for air interface, mobility, and next-generation RAN (NG-RAN), as well as the enhancements on SON and minimization of drive tests (MDT).
@@
The development of 6G follows a process similar to that of 5G. In short, the mobile network industry carries out the process from research and development to testing and implementation. Academia also participates in the research and development work, often cooperating with the industry. Standardization and regulations are handled by the international and national standardization bodies and regulatory organizations. Mobile network operators, service providers, and network vendors carry out the deployment and commercialization phases. The role of ITU-R is to build a vision of IMT for 2030 and beyond (i.e., IMT-2030), define performance/service requirements and evaluation criteria, evaluate candidate RITs, and finally select IMT-2030 compliant (i.e., 6G) systems around 2030. In November 2022, ITU-R published its first technological report on IMT-2030, i.e., Report ITU-R M.2516-0, titled ”Future technology trends of terrestrial IMT systems toward 2030 and beyond”. The report focused on emerging services and applications, technology trends and enablers, and enhanced radio interface and network technologies. In November 2023, ITU-R’s visioning phase culminated in the publication of Recommendation ITU-R M.2160-0, titled ”Framework and overall objectives of the future development of IMT for 2030 and beyond”. This document discussed the trends, usage scenarios, capabilities, and ongoing development of IMT-2030.
The next phase is to define the performance/service requirements and evaluation criteria/methodology of IMT-2030 between 2024 and 2026. The final phase is to evaluate and select compliant RITs and complete IMT-2030 specifications in the time frame of 2027-2030. ITU-R also defines the frequency bands for IMT-2030 through World Radio Conferences (WRCs), organized every three to four years. While WRC 2023 paved the way toward the candidate spectrum bands for IMT-2030 (i.e., 4.4–4.8 GHz, 7.1–8.4 GHz, and 14.8–15.3 GHz), final allocations will be made at the future WRCs, especially at the WRC 2027. 3GPP also plays a key role in the standardization of 6G. When 5G evolution approaches its limits, the standardization efforts start to shift toward 6G. The 6G standardization process is estimated to begin in 2025, Release 20 (7/2025–12/2026) being the first specification to study 6G and Release 21 (1/2027–6/2028) the first official 6G standard. The worldwide commercialization phase of 6G networks is anticipated to start around 2030. The estimated timeline for the development process of 6G is summarized in Figure 3.
@@
The research on 6G was initiated in 2018. This section highlights the primary research steps taken worldwide since then. Many different types of activities are introduced, including initiatives, projects, collaborations, funding, experiments, events, and publications. For clarity, the activities are introduced in a chronological order. Further details on the following 6G activities can be found in [19], where a comprehensive list of 6G research progress is provided. The main benchmarks of 6G research are summarized in Figure 4.
@@ I-D1 Activities in 2018
The first research activities on 6G were conducted in 2018 in terms of projects, events, and publications. The first large-scale 6G project, with a total volume of 249 million euros, was launched in April in Finland by a consortium of the Finnish leading academic and industrial organizations. The 6G Flagship Program is an eight-year research initiative that aims to develop and test 6G enabling technologies. In September, the first annual 6G special issue, entitled ”6G: What is Next?”, was published by the Institute of Electrical and Electronics Engineers (IEEE) Vehicular Technology Magazine. In this special issue, the first 6G vision paper was published under the title ”6G Vision and Requirements: Is There Any Need for Beyond 5G?[20]. The paper contemplated the need for 6G, with a positive conclusion by reviewing each generation’s strengths and deficiencies. In November, another 6G vision paper was published by the Finnish 6G Flagship Program, sketching its tentative vision for 6G and emphasizing the importance of interdisciplinary research in wireless communications, computer engineering, electronics, and material science [21]. In December, Carleton University organized the first workshop series on 6G, covering physical (PHY) layer technologies, network architectures, and AI/ML methods.
@@ I-D2 Activities in 2019
The 6G research activities began to spread around the world in 2019. In January, South Korean LG Electronics opened a center for 6G research at the Korean Advanced Institute of Science and Technology (KAIST). Samsung also established a new 6G research unit called the Advanced Communications Research Center. In March, the Federal Communications Commission announced that it will open the spectrum from 95 GHz to 3 THz for experimental usage and 6G research in the United States (US). In April, the UK launched a five-year project to study optical wireless technologies for 6G and beyond. China Mobile and Tsinghua University reported their cooperation to study 6G technologies in May. In June, SK Telecom partnered with Ericsson and Nokia to conduct 6G research. South Korea and Finland reported joint cooperation in 6G development.
In August, Keysight Technologies joined the Finnish 6G Flagship Program, strengthening its testing and experimental capabilities. Huawei announced that it had already begun to study 6G at its Canadian research center in Ottawa. In November, China’s Ministry of Science and Technology stated that it is going to establish two national 6G groups. One will promote 6G research and development work, consisting of governmental departments, while the other will be formed by academic and industrial organizations, focusing on the technical part of 6G. Rohde &\&& Schwarz reported collaboration with a group of German research institutions to study communications at THz frequencies. The Indian company Tech Mahindra and Finnish government-led Business Finland agency signed a Memorandum of Understanding (MoU) for the joint exploration of 5G/6G, with an agreement to establish an innovation lab in Finland.
A handful of 6G events were arranged around the world in 2019. The first 6G conference was held in March, i.e., the annual 6G Wireless Summit organized by the Finnish 6G Flagship Program. As an outcome of the Summit, the first 6G white paper, ”Key Drivers and Research Challenges for 6G Ubiquitous Wireless Intelligence[22], was published. A few 6G panels/workshops were organized at major conferences. Different research groups around the world began to propose their own visions of what 6G might be in the 2030s. These 6G vision papers discussed many aspects of 6G, such as key trends, main requirements, system architectures, potential technologies, possible applications, opportunities, and challenges. In September, IEEE Vehicular Technology Magazine released its second special issue on 6G.
@@ I-D3 Activities in 2020
6G research continued to spread around the world in 2020. In January, Japan’s government organized a 6G panel with academic and industrial representatives to plan a comprehensive strategy for Japan’s 6G research. In February, the ITU-R Working Party 5D (WP 5D), in charge of IMT systems, decided to begin the visioning work toward IMT-2030. In May, the Chinese technology vendor ZTE and mobile network operator China Unicom agreed to collaborate on 6G research, including the joint development, verification, and trials of 6G enabling technologies. Later in May, the Alliance for Telecommunications Industry Solutions (ATIS) released a call to action to advance the US toward a global leadership in 6G and beyond. The aim was to build close collaboration between the government, academia, and industry to provide a comprehensive path from research and development to standardization and commercialization. In June, Finland’s 6G Flagship Program published eleven 6G white papers, providing a comprehensive study on different aspects of 6G [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Later in autumn, the 6G Flagship Program released an online webinar series, in which each white paper was introduced and thoroughly discussed.
In August, the South Korean government announced that it is planning to allocate 169 million dollars for the development of 6G technologies during the period from 2021 to 2026, and then start a pilot project for 6G services, with five focus areas: immersive content, digital healthcare, autonomous vehicles, smart factories, and smart cities. In October, the Next G Alliance was established to promote 6G in North America. The founding members included numerous major players in the mobile industry. The target of the Next G Alliance is to ensure that North America is a global leader in the 6G development. ITU-R WP 5D released a Liaison statement, inviting external organizations to contribute to the IMT-2030 visioning work by providing views on different technological trends and aspects of 6G. In November, China launched the first experimental ”6G” satellite to test THz communication in space. The University of Surrey opened a 6G Innovation Center and published a white paper, introducing its 6G vision [34].
In 2020, a few 6G conferences were organized virtually. The 6G theme spread to all major telecommunication conferences in the forms of keynote speeches, plenary talks, panels, workshops, and forums. Many major technology companies published white papers to reveal their initial 6G visions. IEEE Vehicular Technology Magazine and IEEE Access released special issues on the 6G-specific topics. Numerous 6G vision and survey papers were also published in 2020.
@@ I-D4 Activities in 2021
At the beginning of 2021, many European Union (EU)-funded 6G projects were launched, including Hexa-X, RISE-6G, 6G BRAINS, and DEDICAT 6G. The Hexa-X initiative is the EU’s 6G flagship project that aims to develop 6G technology enablers to connect human, physical, and digital worlds. RISE-6G studies smart and sustainable wireless environments using RISs, focusing on the propagation modeling, investigation of fundamental limits, algorithmic designs, prototyping, and trialing. 6G BRAINS targets to develop AI-driven resource allocation solutions by exploiting THz and optical frequencies to enable massive connectivity in industrial environments. The objective of the DEDICAT 6G project is to design smart 6G connectivity mechanisms that support the dynamic distribution of intelligence and enable secure human-centric applications.
@@
In June, Vodafone Germany announced that it will open a new 5G/6G research and development center in Dresden. Finland’s 6G Flagship Program and Japan’s Beyond 5G Promotion Consortium agreed on a collaboration for joint research and development on 6G technologies, aiming to contribute to the regulation and standardization of 6G. Samsung and the University of California, Santa Barbara demonstrated an end-to-end sub-THz link with fully digital beamforming at a 140 GHz carrier frequency at the IEEE International Conference on Communications (ICC) 2021. In July, the University of Texas partnered with a group of American and Korean technology companies to establish a joint 6G research and development center, with a special focus on wireless ML, advanced sensing, innovative networking, new spectrum technologies, and satellite communications. Ericsson and the Massachusetts Institute of Technology reported a collaboration on two 6G-related project, focusing on neuromorphic computing and ”zero-energy” devices. The Chinese technology company OPPO announced the establishment of a research team to study 6G requirements, technologies, and system architectures. OPPO also released a 6G white paper on AI-based intelligent networking. Mobile network operators LG UPlus and KDDI reported that they have signed an MoU to study novel business opportunities for 5G networks and develop 6G enabling technologies.
@@
Numerous, mostly virtual or hybrid, 6G events were arranged in 2021, including conferences, workshops, panels, forums, plenary talks, and keynote speeches. In terms of publications, many IEEE journals and magazines released the 6G-specific special issues. Furthermore, many technology companies, academic organizations, and research coalitions published 6G white papers. In addition, several 6G vision and survey papers were published. A handful of the first 6G books were also released.

### v3

@@ I-B7 RELEASE 19: 5G-ADVANCED EVOLUTION
Release 19 (1/2024–6/2025) will be the first evolution of 5G-Advanced. 3GPP agreed on its content in December 2023, and the work began in the beginning of 2024. Release 19 continues the 5G evolution toward 6G, and can be seen largely as the preparation studies toward 6G, while still continuing the development based on commercial deployment needs. One focus will be on developing further use cases, such as the NTN evolution for NR and IoT, femto cells, wireless access/backhaul, low-power wake-up signals, ambient IoT, and multi-hop relays. Another focus area is to further improve the eMBB experience, such as extended reality (XR), mobility, MIMO operation, and duplex. Some interesting starting studies toward 6G technologies include channel modeling for 7-24GHz, which is one of the important new 6G bands in the 6G roadmap and channel modeling for sensing use cases. In addition, improvements of the other key performance indicators will continue in terms of network energy saving enhancements, AI/ML for air interface, and next-generation RAN (NG-RAN), as well as the enhancements on SON and minimization of drive tests (MDT).
@@
In this section, the generic development process of 6G is discussed, covering the big picture and the roles of ITU-R and 3GPP. Generally, ITU-R and 3GPP are vital consortia in 6G standardization, gathering together a large number of stakeholders to develop consistent, interoperable, and global 6G standards. The estimated timeline for the development process of 6G is summarized in Figure 4.
I-C1 BIG PICTURE
The development of 6G follows a process similar to that of 5G. In short, the mobile network industry carries out the process from research and development to testing and implementation. Academia also participates in the research and development work, often cooperating with industry. Standardization and regulations are handled by the international and national standardization bodies and regulatory organizations. In these consortia, there are a vast variety of cooperating stakeholders involved, ranging from companies and universities to alliances and associations. Mobile network operators, service providers, and network vendors carry out the deployment and commercialization phases.
I-C2 ROLE OF ITU-R
ITU acts as a specialized agency of the United Nations for information and communication technologies (ICT) [17]. Its slogan ”committed to connecting the world” reflects the aim of advancing global connectivity and access to digital technologies [17]. In ITU, there are 194 member states and over 1000 companies, universities, and organizations [17]. ITU consists of three sectors, i.e., Radiocommunication (ITU-R), Standardization (ITU-T), and Development (ITU-D) [17]. ITU-R is responsible for the worldwide management of radiocommunication and satellite systems. ITU-T standardizes international ICT systems through its Study Groups full of experts around the world. ITU-D aims to close the digital divide by providing digital services for the least developed countries. From these three sectors, ITU-R plays a key role in 6G standardization. Specifically, the role of ITU-R includes three main phases [18]: phase 1) build a vision of IMT-2030 (2020–2023), phase 2) define performance/service requirements and evaluation methodology/criteria (2024–2026), and phase 3) evaluate candidate RITs and select the IMT-2030 compliant (i.e., 6G) ones by the end of 2030 (2027–2030).
@@ I-C2 ROLE OF ITU-R
ITU-R began its phase 1 visioning work of IMT-2030 in early 2020 and completed it by the end of 2023. The developed IMT-2030 vision was based on two ITU-R documents. In November 2022, ITU-R published its first technological report on IMT-2030, i.e., Report ITU-R M.2516-0, titled ”Future technology trends of terrestrial IMT systems toward 2030 and beyond[19]. The report focused on emerging services and applications, technology trends and enablers, and enhanced radio interface and network technologies. In November 2023, the visioning phase culminated in the publication of Recommendation ITU-R M.2160-0, titled ”Framework and overall objectives of the future development of IMT for 2030 and beyond[20]. This document discussed the trends, usage scenarios, capabilities, and ongoing development of IMT-2030.
The phase 2 of the IMT-2030 development process began in the beginning of 2024 and is expected to be finalized by the end of 2026. Phase 2 is two-fold. First, ITU-R defines the performance and service requirements that IMT-2030 systems need to meet. These requirements will be presented in a report on ”Technical Performance Requirements”. Second, ITU-R develops a comprehensive framework for the evaluation of candidate IMT-2030 systems by defining the evaluation criteria and methodology. This evaluation framework will be introduced in a report on ”Evaluation Methodology”. In phase 2, there will be also a third report, ”Submission Template”, for the candidate RIT proposals.
Phase 3 is planned to begin in early 2027 and be completed by the end of 2030. In phase 3, ITU-R evaluates candidate RITs, developed by 3GPP and other organizations, against the IMT-2030 requirements, relying on the evaluation framework defined in phase 2. Finally, ITU-R selects and approves the RITs that meet the IMT-2030 requirements. These results will be reported in an ITU-R Recommendation document, with a working title of ”Detailed specifications on the terrestrial radio interfaces of international mobile telecommunications-2030 (IMT-2030)” [18]. This document will finalize the work on IMT-2030 systems. The latest details on the IMT-2030 development can be found in [18].
ITU-R plays also another key role in the 6G development process by defining the frequency bands for IMT-2030 through World Radio Conferences (WRCs), organized around every four years. At WRC 2023, ITU-R already began the work toward the candidate spectrum bands for IMT-2030. In Resolution 256, ITU-R defined three frequency ranges from the mid-band spectrum to be studied for IMT-2030 [21], i.e., 4.4–4.8 GHz, 7.125–8.4 GHz, and 14.8–15.35 GHz. These spectrum bands are particularly important for achieving wide coverage since they have favorable propagation characteristics. The plan is to conduct a thorough study of these bands so that they can be approved at WRC 2027.
In addition to the mid-band frequencies, ITU-R defined five spectrum bands from above 100 GHz frequencies for preliminary studies toward possible approval at WRC 2031, i.e., 102–109.5 GHz, 151.5–164 GHz, 167–174.8 GHz, 209–226 GHz and 252–275 GHz, as presented in Resolution 255 [21]. The main promise of these frequency bands is to enable massive network capacity and ultra-broadband communication links in dense hotspot areas. Other potential use cases include ISAC, D2D, and wireless backhaul/fronthaul. For the feasibility of mobile communications at THz frequencies, ITU-R has published a report on the topic in May 2024, i.e., Report ITU-R M.2541-0, titled ”Technical feasibility of IMT in bands above 100 GHz[22]. The conclusion of the report was that it is technically feasible for the IMT-2030 systems to operate at frequencies above 100 GHz, provided that the corresponding challenges are properly resolved. While ITU-R identifies high path, atmospheric, and blocking losses as the main challenges, potential technical solutions include antenna, semiconductor, material, and MIMO technologies.
I-C3 ROLE OF 3GPP
3GPP is a large consortium that develops international standards for mobile telecommunication networks. Since its establishment in 1998, 3GPP has defined standards for 3G, 4G, and 5G networks. This work will continue for 6G as well. 3GPP comprises seven standard development organizations (SDOs) as organizational partners, 26 associations, alliances, and forums as market representation partners, and over 850 companies and organizations as individual members [23]. Whereas organizational partners have the authority to define/publish 3GPP standards and determine the general policies of 3GPP, market representation partners are able to give market advice and guidance to 3GPP. Individual members have the right to contribute to the 3GPP specifications.
Organizational Partners include seven SDOs around the world [23], i.e., the Alliance for Telecommunications Industry Solutions (ATIS) from North America, the European Telecommunications Standards Institute (ETSI) from Europe, China Communications Standards Association (CCSA) from China, Telecommunications Technology Association (TTA) from South Korea, the Association of Radio Industries and Businesses (ARIB) from Japan, Telecommunication Technology Committee (TTC) from Japan, and Telecommunications Standards Development Society, India (TSDSI) from India. These SDOs play a key role in 3GPP and 6G standardization.
3GPP cellular standards are developed through Releases, each of which comprises a major package of new technological features for mobile networks, covering the core network, radio access, terminal, and service capabilities. Every mobile generation contains a set of Releases, each providing an evolutionary step forward. For example, 3G and its evolution was defined in five Releases (99, 4–7), 4G in seven Releases (8–14), and 5G will include six Releases (15–20). Note that there is some overlapping in Releases when the generations are changing, as the current and new/old generation features are developed at the same time. In each Release, new technological features are studied and defined in Technical Reports (TRs) and Technical Specifications (TSs).
Specifically, TRs and TSs are the outputs of study items and work items, respectively. A study item is an exploratory phase to examine the potential and feasibility of new features and technologies, with no commitment to immediate standardization. A work item is a phase to develop exact specifications for new features and technologies, with the intention of producing formal standards. The creation of TRs/TSs is handled by the Technical Specification Groups (TSGs) and their dedicated Working Groups (WGs). In 3GPP, there are three TSGs [23]: Core Network and Terminals (CT), Radio Access Networks (RAN), and Service and System Aspects (SA). RAN and SA TSGs have six WGs each, whereas CT has four. Overall, TSGs are responsible for preparing, approving, and maintaining TRs/TSs.
When 5G evolution approaches its limits, the standardization efforts begin to shift toward 6G. The 6G standardization process is estimated to begin in 2025. If Release 19 will be completed by mid-2025, as planned, the preparation of Release 20 will begin in July 2025. It is planned that Release 20 will begin the 6G standardization process by being the first Release to study 6G features. Assuming the average 18 months preparation time for a Release, Release 20 is anticipated to be finalized by the end of 2026 and the preparation of Release 21 to begin in early 2027. Release 21 is expected to be the first Release to define 6G technologies. Thus, being the first official 6G standard. Most likely Release 21 will be included in the 3GPP submission to ITU-R as a potential 6G radio interface candidate to meet the IMT-2030 requirements. It remains to be seen if Release 22 will complement the 3GPP submission, similarly as Release 16 complemented Release 15 in the IMT-2020 submission. If Release 21 will be completed by mid-2028, as assumed, the first commercial 6G networks could be launched already around late 2028 to early 2029. Thus, the worldwide commercialization phase of 6G networks could begin around 2029–2030.
@@ I-C3 ROLE OF 3GPP
This section highlights the primary research steps taken worldwide since 2018. Many different types of activities are introduced, including initiatives, projects, collaborations, funding, experiments, events, and publications. For clarity, the activities are introduced in a chronological order. Further details on the 6G activities presented in this section can be found in [24], where a comprehensive list of 6G research progress is provided. The main benchmarks of 6G research are summarized in Figure 5. The identification of the main research activities and benchmarks is based on the extensive search of the 6G literature.
@@ I-D1 ACTIVITIES IN 2018
To the best of our knowledge, the first research activities on 6G, which focused solely on 6G and used the term ”6G”, were conducted in 2018 in terms of projects, events, and publications. The first large-scale 6G project [25], with a total volume of 249 million euros, was launched in April in Finland by a consortium of the Finnish leading academic and industrial organizations. The 6G Flagship Program is an eight-year research initiative that aims to develop and test 6G enabling technologies. In September, the first 6G vision paper was published in the Institute of Electrical and Electronics Engineers (IEEE) Vehicular Technology Magazine [26], under the title ”6G Vision and Requirements: Is There Any Need for Beyond 5G?[27]. The paper contemplated the need for 6G, with a positive conclusion by reviewing the strengths and deficiencies of each generation. In November, another 6G vision paper was published by the Finnish 6G Flagship Program, sketching its tentative vision for 6G and emphasizing the importance of interdisciplinary research in wireless communications, computer engineering, electronics, and material science [28]. In December, Carleton University organized the first workshop series on 6G, covering physical (PHY) layer technologies, network architectures, and AI/ML methods.
@@ I-D2 ACTIVITIES IN 2019
6G research activities began to spread around the world in 2019. In January, South Korean LG Electronics opened a center for 6G research at the Korean Advanced Institute of Science and Technology (KAIST). Samsung also established a new 6G research unit called the Advanced Communications Research Center. In March, the Federal Communications Commission announced that it will open the spectrum from 95 GHz to 3 THz for experimental usage and 6G research in the United States (US). In April, the UK launched a five-year project to study optical wireless technologies for 6G and beyond. China Mobile and Tsinghua University reported their cooperation to study 6G technologies in May. In June, SK Telecom partnered with Ericsson and Nokia to conduct 6G research. South Korea and Finland reported joint cooperation in 6G development.
@@
In August, Keysight Technologies joined the Finnish 6G Flagship Program, strengthening its testing and experimental capabilities. Huawei announced that it had already begun to study 6G at its Canadian research center in Ottawa. In November, the Chinese Ministry of Science and Technology stated that it is going to establish two national 6G groups. One will promote 6G research and development work, consisting of governmental departments, while the other will be formed by academic and industrial organizations, focusing on the technical part of 6G. Rohde &\&& Schwarz reported collaboration with a group of German research institutions to study communications at THz frequencies. The Indian company Tech Mahindra and Finnish government-led Business Finland agency signed a Memorandum of Understanding (MoU) for the joint exploration of 5G/6G, with an agreement to establish an innovation lab in Finland.
A handful of 6G events were arranged around the world in 2019. The first conference, fully dedicated on 6G, was held in March, i.e., the annual 6G Wireless Summit organized by the Finnish 6G Flagship Program [29]. As an outcome of the Summit, the first 6G white paper, ”Key Drivers and Research Challenges for 6G Ubiquitous Wireless Intelligence[30], was published. A few 6G panels/workshops were organized at major conferences. Different research groups around the world began to propose their own visions of what 6G might be in the 2030s [31, 32, 33, 34, 35, 36]. These 6G vision papers discussed many aspects of 6G, such as key trends, main requirements, system architectures, potential technologies, possible applications, opportunities, and challenges. In September, IEEE Vehicular Technology Magazine released a special issue on 6G.
@@ I-D3 ACTIVITIES IN 2020
6G research continued to spread around the world in 2020. In January, the government of Japan organized a 6G panel with academic and industrial representatives to plan a comprehensive strategy for the Japanese 6G research. In February, the ITU-R Working Party 5D (WP 5D), in charge of IMT systems, decided to begin the visioning work toward IMT-2030. In May, the Chinese technology vendor ZTE and mobile network operator China Unicom agreed to collaborate on 6G research, including the joint development, verification, and trials of 6G enabling technologies. Later in May, the Alliance for Telecommunications Industry Solutions (ATIS) released a call to action to advance the US toward a global leadership in 6G and beyond. The aim was to build close collaboration between the government, academia, and industry to provide a comprehensive path from research and development to standardization and commercialization. In June, the Finnish 6G Flagship Program published eleven 6G white papers, providing a comprehensive study on different aspects of 6G [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Later in autumn, the 6G Flagship Program released an online webinar series, in which each white paper was introduced and thoroughly discussed.
In July, Samsung published a white paper, revealing its 6G vision on the next hyper-connected experience for all [48]. In August, the South Korean government announced that it is planning to allocate 169 million dollars for the development of 6G technologies during the period from 2021 to 2026, and then start a pilot project for 6G services, with five focus areas: immersive content, digital healthcare, autonomous vehicles, smart factories, and smart cities. In September, Nokia released a white paper on its 6G vision, merging digital, physical, and human worlds [49]. In October, the Next G Alliance was established to promote 6G in North America [50]. The founding members included numerous major players in the mobile industry. The target of the Next G Alliance is to ensure that North America is a global leader in the 6G development. ITU-R WP 5D released a Liaison statement, inviting external organizations to contribute to the IMT-2030 visioning work by providing views on different technological trends and aspects of 6G. In November, China launched the first experimental ”6G” satellite to test THz communication in space. The University of Surrey opened a 6G Innovation Center and published a white paper, introducing its 6G vision [51].
In 2020, a few 6G conferences were organized virtually. The 6G theme spread to all major telecommunication conferences in the forms of keynote speeches, plenary talks, panels, workshops, and forums. Many major technology companies published white papers to reveal their initial 6G visions. IEEE Vehicular Technology Magazine and IEEE Access released special issues on the 6G-specific topics. Numerous 6G visions [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69] and one survey [70] were also published in 2020.
@@ I-D4 ACTIVITIES IN 2021
In March, ITU-R WP 5D established the 6G Vision Group and started the development work toward ”IMT-2030 Framework”. The one6G association was established to promote 6G development by bringing together industrial and academic players around the world [72]. In April, the German government announced that it will fund German 6G research into key technologies and innovative products with 700 million euros by 2025. LG reported a recent collaboration with Keysight Technologies and KAIST to develop 6G enabling technologies, especially for THz communications. A few months later, LG announced a successful demonstration of wireless THz transmission over a 100 meter long outdoor link. Furthermore, LG demonstrated 6G RF front-end technology, jointly developed with Fraunhofer Research Institute, for sub-THz frequencies using the Keysight testbed equipment at the Korean Science and Technology Exhibition 2021.
In June, Vodafone Germany announced that it will open a new 5G/6G research and development center in Dresden. The Finnish 6G Flagship Program and Japanese Beyond 5G Promotion Consortium agreed on a collaboration for joint research and development on 6G technologies, aiming to contribute to the regulation and standardization of 6G. Samsung and the University of California, Santa Barbara demonstrated an end-to-end sub-THz link with fully digital beamforming at a 140 GHz carrier frequency at the IEEE International Conference on Communications (ICC) 2021. In July, the University of Texas partnered with a group of American and Korean technology companies to establish a joint 6G research and development center, with a special focus on wireless ML, advanced sensing, innovative networking, new spectrum technologies, and satellite communications. Ericsson and the Massachusetts Institute of Technology reported a collaboration on two 6G-related project, focusing on neuromorphic computing and ”zero-energy” devices. The Chinese technology company OPPO announced the establishment of a research team to study 6G requirements, technologies, and system architectures. OPPO also released a 6G white paper on AI-based intelligent networking [73]. Mobile network operators LG UPlus and KDDI reported that they have signed an MoU to study novel business opportunities for 5G networks and develop 6G enabling technologies.
@@
Numerous, mostly virtual or hybrid, 6G events were arranged in 2021, including conferences, workshops, panels, forums, plenary talks, and keynote speeches. In terms of publications, many IEEE journals and magazines released the 6G-specific special issues. Furthermore, many technology companies, academic organizations, and research coalitions published 6G white papers. In addition, several 6G visions [77, 78, 79, 80, 76] and surveys [81, 82, 83, 84, 85] were published. A handful of 6G books were also released.